OpenStudy (anonymous):

FAN AND MEDAL Y"ALL!! How many milliliters of magnesium metal, with a density of 1.74 g/mL, are needed to produce 102.3 grams of magnesium oxide in the single-replacement reaction below? Show all steps of your calculation as well as the final answer Mg + H2O → MgO + H2

OpenStudy (aaronq):

So, first step is to balance the equation? is it balanced? Next, you need to work "backwards", to find out how much you need, you first need to find out how much you produced in moles. \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (anonymous):

How do I find the amount of moles in this equation? I'm so bad a Chem :(

OpenStudy (aaronq):

Use the mass given (102.3 grams of magnesium oxide) and the molar mass of the compound. The molar mass can be found by adding the molar masses of the atoms in the formula (or easier by googling it).

OpenStudy (anonymous):

So Google "molar mass of Mg + H2O → MgO + H2"

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

no, there's no such thing. A molar mass is specific for a compound. In this problem, you need to find the moles of one of your products (MgO, magnesium oxide). So you need the molar mass of MgO only.

OpenStudy (anonymous):

Okay hold on

OpenStudy (anonymous):

Okay I got this 40.3044 g/mol

OpenStudy (aaronq):

great. now find the moles of MgO (use the equation i posted above).

OpenStudy (anonymous):

the mass/molar mass?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

I'm tryna figure out the mass, feel like an idiot

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

the mass is given in the question, 102.3 grams

OpenStudy (anonymous):

102.3/ 40.3044 = 2.53818442

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@abb0t help?

OpenStudy (abb0t):

Looks as though someone is already helping. Bye.

OpenStudy (aaronq):

great. Now you need to relate the moles of product to the moles of reactant (they're asking for Mg). To do this, you use a ratio of the moles of each compound and their stoichiometric coefficients. \(\sf \dfrac{moles~of~MgO}{MgO's ~coefficient}=\dfrac{moles~of~Mg}{Mg's ~coefficient}\) plug in your values and find the moles of Mg. (this step is actually sort of redundant but it applies to all problems, so do it anyway).

OpenStudy (anonymous):

Are their coefficients 1?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yes. that's why it's redundant (for this specific exercise). Now we know the moles of Mg. We need to find the mass of Mg with same formula as before - but now with the molar mass of Mg.

OpenStudy (anonymous):

24.3050 g/mol

OpenStudy (aaronq):

okay, find the mass of Mg.

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

nope. use: \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (anonymous):

I'm trying to find mass tho?

OpenStudy (anonymous):

SOOOOO lost

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yeah, just rearrange the formula (it's just algebra). \(\sf mass=Molar~mass*moles\)

OpenStudy (anonymous):

How do I know how many moles are in the formula?

OpenStudy (aaronq):

you found earlier (with the ratio), they're the same as the moles of MgO (because their coefficients are both 1)

OpenStudy (anonymous):

2.53818442 * 24.3050 = 61.69

OpenStudy (aaronq):

good stuff. 61.69 grams now use the density to find the volume (mL) the formula for density is: \(\sf density=\dfrac{mass}{Volume}\)

OpenStudy (anonymous):

How do we know the volume?

OpenStudy (aaronq):

you solve for it by using the mass and the density (this is given in the question)

OpenStudy (anonymous):

So, 1.74 * 61.69 = 107.3406

OpenStudy (aaronq):

nope. check your algebra

OpenStudy (anonymous):

the density in the question is 1.74 tho...?

OpenStudy (aaronq):

right, but the way you solved is wrong (algebraically). the formula is: \(\rho =\dfrac{m}{V}\) when you solve for V, it rearranges to \(V=\dfrac{m}{\rho}\)

OpenStudy (anonymous):

okay so 61.69 / 1.74 = 35.45

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

That is my answer?

OpenStudy (aaronq):

yep. you should be writing the units always to know whats going on

OpenStudy (anonymous):

Final answer to the question is 35.45 grams then?

OpenStudy (anonymous):

@aaronq I was just missing the word grams?

OpenStudy (aaronq):

nope, re-read the question. It's asking you for the mL (volume) of Mg.